OpenStudy (anonymous):

dy/dx-y=ex/x given that y(e)=0

OpenStudy (kainui):

The trick is to multiply this thing by a function to help you undo the product rule. \[y' -y = \frac{e^x}{x} \] Multiply by some function, for some reason people really like to do this with this greek letter \(\mu\) but it's just some function of x. \[\mu y' -\mu y = \mu\frac{e^x}{x} \] Now this is where we undo the product rule, by replacing \(-\mu = \mu'\) \[\mu y' +\mu' y = \mu\frac{e^x}{x} \]\[(\mu y)' = \mu\frac{e^x}{x} \] Since we made up \(\mu\) we solve this new, simpler differential equation we made up and pick the simplest solution, \(-\mu = \mu'\) in other words you can throw away the constant of integration and then plug that in and keep solving for y.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

\[compare ~with~\frac{ dy }{ dx }+Py=Q,where~P~and~Q~are~functions~of~x.\] P=-1 \[Integrating ~factor=e ^{\int\limits Pdx}=e ^{\int\limits -1 dx}=e ^{-x}\] complete solution is \[y.e ^{-x}=\int\limits \frac{ e^x }{ x }e ^{-x}dx+c=\ln \left| x \right|+c\] when x=e y=0 \[0=\ln e+c,0=1+c,c=-1\] ?